OpenStudy (anonymous):

need help simplifying X+7 6 ------ + ------ 2x+12 x^2-36

OpenStudy (anonymous):

so far ive got x+7 6 ---- + ---- 2(x+6) (x+6)(x-6)

OpenStudy (kkutie7):

try finding a common denominator to start.

HanAkoSolo (jamierox4ev3r):

you could also try cross multiplying because it looks like a proportion

OpenStudy (anonymous):

i need the denominators to be the same

OpenStudy (luigi0210):

@tyler123456789 try using @Kkutie7 's way

OpenStudy (anonymous):

how do i find the common denominator

OpenStudy (luigi0210):

Factor

OpenStudy (anonymous):

so.... multiply x-6 to both numerator and denominator on the left side and 2 on the left?

OpenStudy (luigi0210):

Yup, show me what you get

OpenStudy (jdoe0001):

let's take a peek at the denominators 2x+12 can you get a common factor off that?

OpenStudy (jdoe0001):

just that one expression

OpenStudy (luigi0210):

@jdoe0001 he already factored them

OpenStudy (jdoe0001):

ohh, ok, what about \(\bf x^2-36\) keep in mind that \(\bf (a-b)(a+b) = (a^2-b^2)\)

OpenStudy (luigi0210):

@jdoe0001 all under his question

OpenStudy (anonymous):

im confused so when i multiply both i get x+7(x-6) 6(2) --------- + -------- 2(x+6)(x-6) 2(x+6)(x-6)

OpenStudy (jdoe0001):

\(\bf \cfrac{x+7}{2x+12}+\cfrac{6}{x^2-36}\\ \cfrac{x+7}{2(x+6)}+\cfrac{6}{(x+6)(x-6)}\)

OpenStudy (jdoe0001):

the LCD will be really the 2nd denominator

OpenStudy (luigi0210):

Good now multiply the top and continue with the needed operation

OpenStudy (jdoe0001):

\(\bf \cfrac{x+7}{2(x+6)}+\cfrac{6}{(x+6)(x-6)} \implies \cfrac{\square?+\square?}{(x+6)(x-6)}\)

OpenStudy (anonymous):

x^2 + x - 30 ------------ 2(x+6)(x-6)

OpenStudy (jdoe0001):

maybe I needed the 2 as well :/

OpenStudy (luigi0210):

@tyler123456789 there you, that's it

OpenStudy (luigi0210):

Now factor that.

OpenStudy (anonymous):

i dont understand, how?

OpenStudy (anonymous):

nothing sums to 1 and multiplies to 30

OpenStudy (anonymous):

oh wait (x+6)(x-5) --------- 2(x+6)(x-6)

OpenStudy (luigi0210):

\[\frac{(x+6)(x-5)}{2(x+6)(x+6)}\]

OpenStudy (luigi0210):

*(x-6) And correct, now cancel common factors

OpenStudy (anonymous):

si i get x-5 ----- 2x-12

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

thank you so much

OpenStudy (luigi0210):

No problem :)

OpenStudy (anonymous):

new fan